# Gravity as a topological gauge theory

### v1

@@ 3.2 Characteristic poylynomials for $G=SO(4,2)/\{\pm I\}$ and associated action
and after fixing $\alpha=\dfrac{k}{2}\beta$, one gets for the metric:
$g(X,Y)=(\varpi_{\mathfrak{m}}^{*}K_{\mathfrak{m}})(X,Y)=\dfrac{8k}{\ell\ell^{\prime}}\,\eta\bigl{(}\beta(X),\beta(Y)\bigr{)}.$
If one considers the usual covariant derivative $D_{\mathfrak{so}}=(\partial_{\mu}+\dfrac{1}{4}A_{ab,\mu}\gamma^{a}\gamma^{b})dx^{\mu}$, the matter action similar to (2) is then given by :
$S_{M}[A,\beta]$ $=\dfrac{(\ell\ell^{\prime})^{2}}{384(2\ell^{\prime}+k\ell)k}\int_{\mathscr{M}}h^{\varepsilon}\bigl{(}i\psi,K_{\mathfrak{m},ab}\gamma^{a}\varpi_{\mathfrak{m}}^{b}\wedge*D_{\mathfrak{so}}\psi\bigl{)}=ReS_{D}$ (7)
A more natural matter action $S^{\prime}_{M}$ with covariant derivative $D_{0}=d+\Phi_{\varepsilon}(\varpi_{0})=d+\Phi_{\varepsilon}(A-\lambda\mathds{1}_{4})=(\partial_{\mu}+\dfrac{1}{4}A_{ab,\mu}\gamma^{a}\gamma^{b}-3\lambda_{\mu})dx^{\mu}$ taking into account the full action of the Weyl algebra $\mathfrak{g}_{0}=\mathfrak{co}(3,2)$-valued part of the connection $\varpi_{0}$ is:
$S^{\prime}_{M}[A,\beta,\lambda]$ $=\dfrac{(\ell\ell^{\prime})^{2}}{384(2\ell^{\prime}+k\ell)k}\int_{\mathscr{M}}h^{\varepsilon}\bigl{(}i\psi,K_{\mathfrak{m},ab}\gamma^{a}\varpi_{\mathfrak{m}}^{b}\wedge*D_{0}\psi\bigl{)}$ (8)
$=\overbrace{S_{M}}^{ReS_{D}}-\dfrac{(\ell\ell^{\prime})^{2}}{128(2\ell^{\prime}+k\ell)k}\int_{\mathscr{M}}h^{\varepsilon}\bigl{(}i\psi,K_{\mathfrak{m},ab}\gamma^{a}\varpi_{\mathfrak{m}}^{b}\wedge*(\lambda\psi)\bigl{)}.$ (9)
The study of the full action $S^{\prime}_{T}=S_{G}+S^{\prime}_{M}$ is left for a future work. In the following we will restrict ourselves to the study of the action $S_{T}=S_{G}+S_{M}$. For $k\neq 0$ the equations of motion for $\beta$, $A$ and $\lambda$ yield (neglecting boundary terms) :
$\dfrac{\delta\mathscr{L}_{G}[A,\beta,\lambda]}{\delta\beta^{c}}$ $=\dfrac{ek}{4\pi^{2}\ell\ell^{\prime}}(R^{ab}+\dfrac{2k}{\ell\ell^{\prime}}\beta^{a}\wedge\beta^{b})\wedge\beta^{d}\varepsilon_{abcd}+\dfrac{rk}{2\pi^{2}\ell\ell^{\prime}}(2R^{ab}\wedge\beta_{b}\delta_{ac}+T_{c}\wedge\lambda-\beta_{c}\wedge d\lambda)$
@@
$\dfrac{\delta\mathscr{L}_{G}[A,\beta,\lambda]}{\delta A^{ab}}$ $=\dfrac{ek}{4\pi^{2}\ell\ell^{\prime}}T^{c}\wedge\beta^{d}\varepsilon_{abcd}+\dfrac{rk}{2\pi^{2}\ell\ell^{\prime}}(2T_{a}+\lambda\wedge\beta_{a})\wedge\beta_{b}=\dfrac{1}{2}\mathfrak{s}_{ab}=-\dfrac{\delta\mathscr{L}_{M}[A,\beta,\lambda]}{\delta A^{ab}}$ (11)
@@
Equation (12) implies either $r=0$ or $T=0$ since the definition of a Cartan connection requires $\beta$ to be an isomorphism. For $r=0$, one then recovers the usual equations of motion associated to Einstein-Cartan gravity, with the torsion $T$ linked to the spin density of matter. In the other case $T=0$, the equations of motion become:
$\dfrac{ek}{4\pi^{2}\ell\ell^{\prime}}(R^{ab}+\dfrac{2k}{\ell\ell^{\prime}}\beta^{a}\wedge\beta^{b})\wedge\beta^{d}\varepsilon_{abcd}+\dfrac{rk}{2\pi^{2}\ell\ell^{\prime}}(2R^{ab}\wedge\beta_{b}\delta_{ac}-\beta_{c}\wedge d\lambda)$ $=\tau_{c}$ (13)
$\dfrac{rk}{2\pi^{2}\ell\ell^{\prime}}\lambda\wedge\beta_{a}\wedge\beta_{b}$ $=\dfrac{1}{2}\mathfrak{s}_{ab}$ (14)
@@
$\lambda_{c}=\lambda_{\mu}\beta^{-1\,\mu}_{~{}~{}\,\,\;\;\;c}=\dfrac{\pi^{2}\ell\ell^{\prime}}{12rk}\mathfrak{s}^{ab}_{~{}~{}abc}\ .$ (16)
Using the relation $\gamma=\dfrac{2\alpha_{1}}{2\alpha_{2}-\alpha_{5}}$, given in [14, p13], we can identify $\alpha_{5}=-\alpha_{2}=-\dfrac{rk}{4\pi^{2}\ell\ell^{\prime}}=-\dfrac{2\alpha_{1}}{3\gamma}$, with $\gamma$ the Barbero-Immirzi parameter. Provided we set $e=\dfrac{8\pi^{2}\alpha_{1}\ell\ell^{\prime}}{k}=\dfrac{3\pi}{2\Lambda_{0}G}$ and $\ell\ell^{\prime}=-\dfrac{6k}{\Lambda_{0}}$ where $\alpha_{1}=-\dfrac{1}{16\pi G}$, and applying the Hodge star *** to the first equation and replacing $\lambda$ according to (16) gives us the Einstein equations modified by the Barbero-Immirzi part of the action with bare cosmological constant $\Lambda_{0}$ and with matter sources given by the energy-momentum tensor and a particular combination of derivatives of the spin density of matter (the latter resulting from the presence of the dilation 1-form $\lambda$ of the connection):
$G_{kc}+\dfrac{2}{3\gamma}R^{ab}_{~{}~{}rs}\varepsilon^{rs}_{~{}~{}bk}\delta_{ac}+\Lambda_{0}\eta_{kc}=-\dfrac{1}{2\alpha_{1}}(\tau_{ck}-\dfrac{1}{48}\partial_{\mu}\mathfrak{s}^{ab}_{~{}~{}ab\nu}\varepsilon^{\mu\nu}_{~{}~{}\;ck}).$ (17)
@@
$G_{kc}+\dfrac{2}{3\gamma}R^{ab}_{~{}~{}rs}\varepsilon^{rs}_{~{}~{}bk}\delta_{ac}+\Lambda\eta_{kc}=-\dfrac{1}{2\alpha_{1}}(\tau_{M,ck}-\dfrac{1}{48}\partial_{\mu}\mathfrak{s}^{ab}_{~{}~{}ab\nu}\varepsilon^{\mu\nu}_{~{}~{}\;ck})$ (18)
with $\Lambda=\Lambda_{0}+\Lambda_{\text{vac}}=\overbrace{-\dfrac{6k}{\ell\ell^{\prime}}}^{\Lambda_{0}}\overbrace{-\dfrac{\rho_{\text{vac}}}{2\alpha_{1}}}^{\Lambda_{\text{vac}}}=-\dfrac{6k}{\ell\ell^{\prime}}-\dfrac{2\pi^{2}\ell\ell^{\prime}\rho_{\text{vac}}}{ek}=\Lambda_{0}+8\pi G\rho_{\text{vac}}$ the effective cosmological constant, and $\Lambda_{V}$ the vacuum contribution to the cosmological constant. One can infer that changes in cosmological history of the rate of expansion of the universe could be partly linked to spin-density variations in this way.
@@
In this particular case the fact that Torsion is null implies that the Einstein tensor $G_{kc}$ is symmetric (see [15, p3] for more details). When taking into account that $\eta_{kc}$ is symmetric as well, the anti-symmetricity of $\partial_{\mu}\mathfrak{s}^{ab}_{~{}~{}ab\nu}\varepsilon^{\mu\nu}_{~{}~{}\;ck}$ implies that the energy-momentum tensor $\tau_{ck}$ sports an anti-symmetric part if $\partial_{\mu}\mathfrak{s}^{ab}_{~{}~{}ab\nu}\varepsilon^{\mu\nu}_{~{}~{}\;ck}\neq 0$. This confirms that we can consider a matter Lagrangian that depends on the spin connection $A$ as we explicitly did.
Replacing $e=\dfrac{8\pi^{2}\alpha_{1}\ell\ell^{\prime}}{k}=\dfrac{3\pi}{2\Lambda_{0}G}$ and $\ell\ell^{\prime}=-\dfrac{6k}{\Lambda_{0}}$ in the action and $\dfrac{rk}{4\pi^{2}\ell\ell^{\prime}}=\dfrac{2\alpha_{1}}{3\gamma}$, we obtain:
$S_{G}[\varpi]$ $=\int_{\mathscr{M}}\biggl{(}\overbrace{\alpha_{1}(R^{ab}\wedge\beta^{c}\wedge\beta^{d}\varepsilon_{abcd}+\dfrac{2}{3\gamma}R^{ab}\wedge\beta_{a}\wedge\beta_{b})}^{Holst}+\overbrace{\dfrac{\Lambda_{0}\alpha_{1}}{6}\beta^{a}\wedge\beta^{b}\wedge\beta^{c}\wedge\beta^{d}\varepsilon_{abcd}}^{Cosmological~{}constant}$
$\qquad\qquad+\overbrace{\dfrac{3\alpha_{1}}{2\Lambda_{0}}R^{ab}\wedge R^{cd}\varepsilon_{abcd}}^{Euler}+\overbrace{\underbrace{\dfrac{2\alpha_{1}}{\Lambda_{0}\gamma}}_{\vartheta}R^{ab}\wedge R_{ab}}^{Pontrjagin}\overbrace{-\dfrac{2\alpha_{1}}{3\gamma}\bigl{(}T^{a}\wedge T_{a}-R^{ab}\wedge\beta_{a}\wedge\beta_{b}\bigr{)}}^{Nieh-Yan}$
$\qquad\qquad\overbrace{-\dfrac{4\alpha_{1}}{3\gamma}\eta_{ab}T^{a}\wedge\lambda\wedge\beta^{b}}^{\text{ Dilation-Torsion interaction}}+\overbrace{\underbrace{\dfrac{4\alpha_{1}}{\Lambda_{0}\gamma}}_{-2\vartheta}d\lambda\wedge d\lambda}^{\text{ Dilation kinetic term }}\biggr{)}.$ (19)
@@
$V(\vartheta)$ $=\dfrac{2\alpha_{1}}{3\gamma}R^{ab}\wedge\beta_{a}\wedge\beta_{b}+\overbrace{\dfrac{\Lambda_{0}\alpha_{1}}{6}\beta^{a}\wedge\beta^{b}\wedge\beta^{c}\wedge\beta^{d}\varepsilon_{abcd}}^{Cosmological~{}constant}+\overbrace{\dfrac{3\alpha_{1}}{2\Lambda_{0}}R^{ab}\wedge R^{cd}\varepsilon_{abcd}}^{Euler}$
$\qquad\qquad\overbrace{-\dfrac{2\alpha_{1}}{3\gamma}\bigl{(}T^{a}\wedge T_{a}-R^{ab}\wedge\beta_{a}\wedge\beta_{b}\bigr{)}}^{Nieh-Yan}\overbrace{-\dfrac{4\alpha_{1}}{3\gamma}\eta_{ab}T^{a}\wedge\lambda\wedge\beta^{b}}^{\text{ Dilation-Torsion interaction}}-\overbrace{2\vartheta d\lambda\wedge d\lambda}^{\text{ Dilation kinetic term }}$ (20)
@@
$S_{G}[\varpi]$ $=\int_{\mathscr{M}}\biggl{(}\overbrace{\alpha_{1}(R^{ab}\wedge\beta^{c}\wedge\beta^{d}\varepsilon_{abcd}+\dfrac{2}{3\gamma}R^{ab}\wedge\beta_{a}\wedge\beta_{b})}^{Holst}+\overbrace{\dfrac{\Lambda_{0}\alpha_{1}}{6}\beta^{a}\wedge\beta^{b}\wedge\beta^{c}\wedge\beta^{d}\varepsilon_{abcd}}^{Cosmological~{}constant}$
$\qquad\qquad+\overbrace{\dfrac{3\alpha_{1}}{2\Lambda_{0}}R^{ab}\wedge R^{cd}\varepsilon_{abcd}}^{Euler}+\overbrace{\underbrace{\dfrac{2\alpha_{1}}{\Lambda_{0}\gamma}}_{\vartheta}R^{ab}\wedge R_{ab}}^{Pontrjagin}\overbrace{-\dfrac{2\alpha_{1}}{3\gamma}\bigl{(}T^{a}\wedge T_{a}-R^{ab}\wedge\beta_{a}\wedge\beta_{b}\bigr{)}}^{Nieh-Yan}\biggr{)}\Biggr{)}.$
@@
$\left\{\begin{array}[]{ll}&\lambda=0\text{ and }\alpha=\dfrac{k}{2}\beta\\ &\left\{\begin{array}[]{ll}sign(k\ell\ell^{\prime})=-,\quad\Lambda_{0}=-\dfrac{6k}{\ell\ell^{\prime}}>0&\Leftrightarrow G=SO(4,1)\\ k=0&\Leftrightarrow G=ISO(3,1)\\ sign(k\ell\ell^{\prime})=+,\quad\Lambda_{0}=-\dfrac{6k}{\ell\ell^{\prime}}<0&\Leftrightarrow G=SO(3,2)\end{array}\right.\end{array}\right.$ (26)
Remark.
$\alpha_{2}=0$ and $\lambda=0$ give different actions but the same equations of motion if we neglect the boundary terms.

### v3

@@ 3.2 Characteristic poylynomials for $G=SO(4,2)/\{\pm I\}$ and associated action
and after fixing $\alpha=\dfrac{k}{2}\beta$, one gets for the metrics:
$h(\gamma,\gamma^{\prime})=\kappa K_{\mathfrak{g}}(\gamma,\gamma^{\prime})$ (7)
$g(X,Y)=\zeta(\varpi_{\mathfrak{m}}^{*}h)(X,Y)=\kappa\zeta(\varpi_{\mathfrak{m}}^{*}K_{\mathfrak{m}})(X,Y)=\dfrac{8k\kappa\zeta}{\ell\ell^{\prime}}\,\eta\bigl{(}\beta(X),\beta(Y)\bigr{)}.$ (8)
If one considers the usual covariant derivative $D_{\mathfrak{so}}=(\partial_{\mu}+\dfrac{1}{4}A_{ab,\mu}\gamma^{a}\gamma^{b})dx^{\mu}$, the matter action similar to (3) is then given by  [^9] Where we set $\kappa^{2}\zeta=\tfrac{(\ell\ell^{\prime})^{2}}{192(2\ell^{\prime}+k\ell)k}$ such that the overall factor is arranged to stick with the Dirac action given in [7, p.197].:
$S_{M}[A,\beta]$ $=\int_{\mathscr{M}}h^{\varepsilon}\bigl{(}i\psi,h_{ab}\gamma^{a}\varpi_{\mathfrak{m}}^{b}\wedge*D_{\mathfrak{so}}\psi\bigl{)}=\mathrm{Re}S_{D}\,.$ (9)
Let $C_{i}$ be the $i$-dimensional Clifford algebra. We can extend the isomorphism 2 between $\mathfrak{so}(n-1,1)$ and $\mathfrak{spin}(n-1,1)$ to the morphism $\Phi_{\varepsilon}:\mathfrak{gl}(n,\mathbb{R})\rightarrow C_{2}\oplus C_{0}$. Let $M=M^{rs}e_{r}\bar{e}_{s}\in\mathfrak{gl(n,\mathbb{R)}}$. We define:
$\Phi_{\varepsilon}(M)$ $=M^{rs}\Phi_{\varepsilon}(e_{r}\bar{e}_{s})=\dfrac{M^{rs}}{4}\bar{e}_{a}e_{r}\bar{e}_{s}e_{b}\gamma^{a}\gamma^{b}=\dfrac{M^{rs}}{4}\eta_{ar}\eta_{sb}\gamma^{a}\gamma^{b}$ (10)
$=\dfrac{1}{4}\bigl{(}\sum_{a<b}M_{ab}\gamma^{a}\gamma^{b}+\sum_{a>b}M_{ab}\gamma^{a}\gamma^{b}+\sum_{a}M_{aa}(\gamma^{a})^{2}\bigr{)}$ (11)
$=\sum_{a<b}\dfrac{1}{4}(M_{ab}-M_{ba})\gamma^{a}\gamma^{b}+\dfrac{M_{ab}}{4}\eta^{ab}$ (12)
where we used the relation $(\gamma^{a})^{2}=\left\{\begin{array}[]{lcll}+1\text{ if }a\leq n-1\\ -1\text{ if }a>n-1\end{array}\right.$. A particular case is the isomorphism between elements of the algebra $\mathfrak{g}_{0}=\mathfrak{so}(n-1,1)\oplus\mathbb{R}$ and those of $\mathfrak{spin}(n-1,1)\oplus\mathbb{R}$ which corresponds to $\Phi_{\varepsilon}:\mathfrak{so}(n-1,1)\oplus\mathbb{R}\rightarrow\mathfrak{spin}(n-1,1)\oplus\mathbb{R}$. The action of $\Phi_{\varepsilon}$ on the generators of $\mathfrak{so}(n-1,1)$ and on the identity matrix which generates the dilation part of $\mathfrak{g}$ is given by:
$\Phi_{\varepsilon}(J_{rs}\oplus\mathds{1}_{n})=\dfrac{1}{4}\bar{e}_{a}(J_{rs}\oplus\mathds{1})e_{b}\gamma^{a}\gamma^{b}=\dfrac{1}{4}\bigl{(}(J_{rs})_{ab}\oplus\bar{e}_{a}e_{r}\bar{e}_{s}e_{b}\delta^{rs}\bigr{)}\gamma^{a}\gamma^{b}=\dfrac{1}{2}\eta_{ar}\eta_{bs}\gamma^{a}\gamma^{b}\oplus\dfrac{n-2}{4}.$ (13)
Thus, a more natural matter action $S^{\prime}_{M}$ with covariant derivative $D_{0}=d+\Phi_{\varepsilon}(\varpi_{0})=d+\Phi_{\varepsilon}(A-\lambda\mathds{1}_{4})=(\partial_{\mu}+\dfrac{1}{4}A_{ab,\mu}\gamma^{a}\gamma^{b}-\dfrac{1}{2}\lambda_{\mu})dx^{\mu}$ taking into account the full action of the $\mathfrak{g}_{0}=\mathfrak{co}(3,2)$-valued part of the connection $\varpi_{0}$ is:
$S^{\prime}_{M}[A,\beta,\lambda]$ $=\int_{\mathscr{M}}h^{\varepsilon}\bigl{(}i\psi,h_{ab}\gamma^{a}\varpi_{\mathfrak{m}}^{b}\wedge*D_{0}\psi\bigl{)}$ (14)
$=\overbrace{S_{M}}^{\mathrm{Re}S_{D}}-\dfrac{1}{2}\int_{\mathscr{M}}h^{\varepsilon}\bigl{(}i\psi,h_{ab}\gamma^{a}\varpi_{\mathfrak{m}}^{b}\wedge*(\lambda\psi)\bigl{)}.$ (15)
The study of the full action $S^{\prime}_{T}=S_{G}+S^{\prime}_{M}$ is left for future work. In the following we will restrict ourselves to the study of the action $S_{T}=S_{G}+S_{M}$.
For $k\neq 0$ the equations of motion for $\beta$, $A$ and $\lambda$ yield (neglecting boundary terms):
$\dfrac{\delta\mathscr{L}_{G}[A,\beta,\lambda]}{\delta\beta^{c}}$ $=\dfrac{ek}{4\pi^{2}\ell\ell^{\prime}}(R^{ab}+\dfrac{2k}{\ell\ell^{\prime}}\beta^{a}\wedge\beta^{b})\wedge\beta^{d}\varepsilon_{abcd}+\dfrac{k}{2\pi^{2}\ell\ell^{\prime}}(2(r+y)R^{ab}\wedge\beta_{b}\delta_{ac}+rT_{c}\wedge\lambda-r\beta_{c}\wedge d\lambda)$
@@
$\dfrac{\delta\mathscr{L}_{G}[A,\beta,\lambda]}{\delta A^{ab}}$ $=\dfrac{ek}{4\pi^{2}\ell\ell^{\prime}}T^{c}\wedge\beta^{d}\varepsilon_{abcd}+\dfrac{k}{\pi^{2}\ell\ell^{\prime}}\bigl{(}(\tfrac{r}{2}+y)T_{a}+\dfrac{r}{4}\lambda\wedge\beta_{a}\bigr{)}\wedge\beta_{b}=\dfrac{1}{2}\mathfrak{s}_{ab}=-\dfrac{\delta\mathscr{L}_{M}[A,\beta,\lambda]}{\delta A^{ab}}$ (17)
@@
Equation (18) implies either $r=0$ or $T^{a}\wedge\beta_{a}=0$.
For $r=0$, one then recovers the usual equations of motion associated to Einstein-Cartan gravity with a Holst action and bare cosmological constant term, and torsion $T$ linked to the spin density of matter.
In the other case, using the Hodge star operator lets us show that $T^{a}\wedge\beta_{a}=0$ is equivalent to requiring the axial vector part of torsion to vanish ($\tilde{t}^{a}=\dfrac{1}{6}T^{b}_{~{}\mu\nu}\varepsilon^{a~{}\mu\nu}_{~{}b}=0$) as is described in [15, Table 3]. The end result being that the new torsion $T^{a}_{~{}bc}=Z^{a}_{~{}bc}+\delta^{a}_{[b}t^{\vphantom{a}}_{c]}$ can be expressed solely in terms of $20$ independent components instead of the usual $24$. Here $Z^{a}_{~{}bc}$ is a $3$-indices tensor and $t_{c}$ is a $4$-dimensional vector with respectively $16$ and $4$ independent components.
A very particular case satisfying $T^{a}\wedge\beta_{a}=0$ is $T=0$, the equations of motion then become:
$\dfrac{ek}{4\pi^{2}\ell\ell^{\prime}}(R^{ab}+\dfrac{2k}{\ell\ell^{\prime}}\beta^{a}\wedge\beta^{b})\wedge\beta^{d}\varepsilon_{abcd}+\dfrac{k}{2\pi^{2}\ell\ell^{\prime}}\bigl{(}2(r+y)R^{ab}\wedge\beta_{b}\delta_{ac}-r\beta_{c}\wedge d\lambda\bigr{)}$ $=\tau_{c}$ (19)
$\dfrac{rk}{4\pi^{2}\ell\ell^{\prime}}\lambda\wedge\beta_{a}\wedge\beta_{b}$ $=\dfrac{1}{2}\mathfrak{s}_{ab}$ (20)
@@
$\lambda_{c}=\lambda_{\mu}\beta^{-1\,\mu}_{~{}~{}\,\,\;\;\;c}=\dfrac{\pi^{2}\ell\ell^{\prime}}{3rk}\mathfrak{s}^{ab}_{~{}~{}abc}\ .$ (22)
Using the relation $\gamma=\dfrac{2\alpha_{1}}{2\alpha_{2}-\alpha_{5}}$, given in [8, p13], we identify $\gamma=\dfrac{2e}{3r+4y}$, with $\gamma$ the Barbero-Immirzi parameter.
Provided we set $e=\dfrac{4\pi^{2}\alpha_{1}\ell\ell^{\prime}}{k}=\dfrac{3\pi}{2\Lambda_{0}G}$ and $\ell\ell^{\prime}=-\dfrac{6k}{\Lambda_{0}}$ where $\alpha_{1}=-\dfrac{1}{16\pi G}$, applying the Hodge star * to (19) and replacing $\lambda$ according to (22) gives us the Einstein equations modified by the Barbero-Immirzi part of the action with bare cosmological constant $\Lambda_{0}$ and with matter sources given by the energy-momentum tensor and a particular combination of derivatives of the spin density of matter (the latter resulting from the presence of the dilation 1-form $\lambda$ of the connection):
$G_{kc}+(\dfrac{\Lambda_{0}Gr}{6\pi}+\dfrac{1}{2\gamma})R^{ab}_{~{}~{}rs}\varepsilon^{rs}_{~{}~{}bk}\delta_{ac}+\Lambda_{0}\eta_{kc}=-\dfrac{1}{2\alpha_{1}}(\tau_{ck}-\dfrac{1}{12}\partial_{\mu}\mathfrak{s}^{ab}_{~{}~{}ab\nu}\varepsilon^{\mu\nu}_{~{}~{}\;ck}).$ (23)
@@
$G_{kc}+(\dfrac{\Lambda_{0}Gr}{6\pi}+\dfrac{1}{2\gamma})R^{ab}_{~{}~{}rs}\varepsilon^{rs}_{~{}~{}bk}\delta_{ac}+\Lambda\eta_{kc}=-\dfrac{1}{2\alpha_{1}}(\tau_{M,ck}-\dfrac{1}{12}\partial_{\mu}\mathfrak{s}^{ab}_{~{}~{}ab\nu}\varepsilon^{\mu\nu}_{~{}~{}\;ck})$ (24)
with $\Lambda=\Lambda_{0}+\Lambda_{\text{vac}}=\overbrace{-\dfrac{6k}{\ell\ell^{\prime}}}^{\Lambda_{0}}\overbrace{-\dfrac{\rho_{\text{vac}}}{2\alpha_{1}}}^{\Lambda_{\text{vac}}}=-\dfrac{6k}{\ell\ell^{\prime}}-\dfrac{2\pi^{2}\ell\ell^{\prime}\rho_{\text{vac}}}{ek}=\Lambda_{0}+8\pi G\rho_{\text{vac}}$ the effective cosmological constant, and $\Lambda_{\text{vac}}$ the vacuum contribution to the cosmological constant.
One can infer that changes in cosmological history of the rate of expansion of the universe could be partly linked to spin-density variations in this way.
If one does not assume $T=0$ the EOM’s instead yield:
$G_{kc}$ $+(\dfrac{\Lambda_{0}Gr}{6\pi}+\dfrac{1}{2\gamma})R^{ab}_{~{}~{}rs}\varepsilon^{rs}_{~{}~{}bk}\delta_{ac}+\Lambda_{0}\eta_{kc}$
$=-\dfrac{1}{2\alpha_{1}}\Bigl{(}\tau_{ck}-\dfrac{1}{24}(2\eta_{cs}\beta^{s}_{\nu}\partial_{\mu}+T_{c,\mu\nu})\bigl{(}\mathfrak{s}^{ab}_{~{}~{}abr}+\dfrac{4(r+2y)}{\pi^{2}}T^{m}_{~{}rm}\bigr{)}\varepsilon^{\mu\nu r}_{~{}~{}~{}\,k}\Bigr{)}$ (25)
$\lambda_{c}$ $=\dfrac{1}{r}\bigl{(}\dfrac{\pi^{2}k}{3\ell\ell^{\prime}}\mathfrak{s}^{ab}_{~{}~{}abc}-8(r+2y)T^{a}_{~{}ca}\bigr{)}$ (26)
$\tilde{t^{a}}$ $=0.$ (27)
@@
In this particular case the fact that Torsion is null implies that the Einstein tensor $G_{kc}$ is symmetric (see [16, p3] for more details). When taking into account that $\eta_{kc}$ is symmetric as well, the anti-symmetricity of $\partial_{\mu}\mathfrak{s}^{ab}_{~{}~{}ab\nu}\varepsilon^{\mu\nu}_{~{}~{}\;ck}$ implies that the energy-momentum tensor $\tau_{ck}$ sports an anti-symmetric part if $\partial_{\mu}\mathfrak{s}^{ab}_{~{}~{}ab\nu}\varepsilon^{\mu\nu}_{~{}~{}\;ck}\neq 0$. This confirms that we can consider a matter Lagrangian that depends on the spin connection $A$ as we explicitly did.
Replacing $e=\dfrac{4\pi^{2}\alpha_{1}\ell\ell^{\prime}}{k}=\dfrac{3\pi}{2\Lambda_{0}G}$; $\ell\ell^{\prime}=-\dfrac{6k}{\Lambda_{0}}$ and $y=\dfrac{3}{4}(\dfrac{\pi}{\Lambda_{0}G\gamma}-r)$ in the action, we obtain:
$S_{G}[\varpi]$ $=\int_{\mathscr{M}}\biggl{(}\dfrac{\alpha_{1}}{2}\Bigl{(}\overbrace{R^{ab}\wedge\beta^{c}\wedge\beta^{d}\varepsilon_{abcd}+(\dfrac{4}{\gamma}-\dfrac{4\Lambda_{0}Gr}{3\pi})R^{ab}\wedge\beta_{a}\wedge\beta_{b}}^{Holst}\overbrace{-\dfrac{\Lambda_{0}}{6}\beta^{a}\wedge\beta^{b}\wedge\beta^{c}\wedge\beta^{d}\varepsilon_{abcd}}^{Cosmological~{}constant}$
$\qquad\qquad\overbrace{-\dfrac{3}{2\Lambda_{0}}R^{ab}\wedge R^{cd}\varepsilon_{abcd}}^{Euler}+\overbrace{\underbrace{(\dfrac{Gr}{\pi}-\dfrac{3}{\Lambda_{0}\gamma})}_{\vartheta}R^{ab}\wedge R_{ab}}^{Pontrjagin}\overbrace{-\dfrac{4\Lambda_{0}Gr}{3\pi}\bigl{(}T^{a}\wedge T_{a}-R^{ab}\wedge\beta_{a}\wedge\beta_{b}\bigr{)}}^{Nieh-Yan}\Bigr{)}$
$\qquad\qquad\overbrace{+\dfrac{\Lambda_{0}r}{12\pi^{2}}\eta_{ab}T^{a}\wedge\lambda\wedge\beta^{b}}^{\text{ Dilation-Torsion interaction}}\overbrace{-\dfrac{r}{4\pi^{2}}d\lambda\wedge d\lambda}^{\text{ Dilation kinetic term }}\biggr{)}.$ (28)
@@
$V(\vartheta)$ $=(\dfrac{2}{\gamma}-\dfrac{2\Lambda_{0}Gr}{3\pi})\alpha_{1}R^{ab}\wedge\beta_{a}\wedge\beta_{b}\overbrace{-\dfrac{\Lambda_{0}\alpha_{1}}{12}\varepsilon_{abcd}\beta^{a}\wedge\beta^{b}\wedge\beta^{c}\wedge\beta^{d}}^{Cosmological~{}constant}\overbrace{-\dfrac{3\alpha_{1}}{4\Lambda_{0}}R^{ab}\wedge R^{cd}\varepsilon_{abcd}}^{Euler}$ (29)
$\qquad\overbrace{-\dfrac{2r\Lambda_{0}G\alpha_{1}}{3\pi}\bigl{(}T^{a}\wedge T_{a}-R^{ab}\wedge\beta_{i}\wedge\beta_{j}\bigr{)}}^{Nieh-Yan}\overbrace{+\dfrac{\Lambda_{0}r}{12\pi^{2}}\eta_{ab}T^{a}\wedge\lambda\wedge\beta^{b}}^{\text{ Dilation-Torsion interaction}}\overbrace{-\dfrac{r}{4\pi^{2}}d\lambda\wedge d\lambda}^{\text{ Dilation kinetic term }}.$ (30)
@@
$S_{G}[\varpi]$ $=\int_{\mathscr{M}}\dfrac{\alpha_{1}}{2}\biggl{(}\overbrace{R^{ab}\wedge\beta^{c}\wedge\beta^{d}\varepsilon_{abcd}+(\dfrac{4}{\gamma}-\dfrac{4\Lambda_{0}Gr}{3\pi})R^{ab}\wedge\beta_{a}\wedge\beta_{b}}^{Holst}\overbrace{-\dfrac{\Lambda_{0}}{6}\beta^{a}\wedge\beta^{b}\wedge\beta^{c}\wedge\beta^{d}\varepsilon_{abcd}}^{Cosmological~{}constant}$
$\qquad\qquad\overbrace{-\dfrac{3}{2\Lambda_{0}}R^{ab}\wedge R^{cd}\varepsilon_{abcd}}^{Euler}+\overbrace{(\dfrac{Gr}{\pi}-\dfrac{3}{\Lambda_{0}\gamma})R^{ab}\wedge R_{ab}}^{Pontrjagin}\overbrace{-\dfrac{4\Lambda_{0}Gr}{3\pi}\bigl{(}T^{a}\wedge T_{a}-R^{ab}\wedge\beta_{a}\wedge\beta_{b}\bigr{)}}^{Nieh-Yan}\biggr{)}.$
@@
$\left\{\begin{array}[]{ll}&\lambda=0\text{ and }\alpha=\dfrac{k}{2}\beta\\ &\left\{\begin{array}[]{ll}sign(k\ell\ell^{\prime})=-,\quad\Lambda_{0}=-\dfrac{6k}{\ell\ell^{\prime}}>0&\Leftrightarrow G=SO(4,1)\\ k=0&\Leftrightarrow G=ISO(3,1)\\ sign(k\ell\ell^{\prime})=+,\quad\Lambda_{0}=-\dfrac{6k}{\ell\ell^{\prime}}<0&\Leftrightarrow G=SO(3,2).\end{array}\right.\end{array}\right.$ (36)